OpenStudy (anonymous):

What polynomial has roots of -6, 1, and 4 ?

OpenStudy (anonymous):

(x+6) * (x-1) * (x-4)

OpenStudy (anonymous):

(x+6) * (x-1) = (x^2 + 5x - 6) (x^2 + 5x - 6) * (x-4) = (x^3 +5x^2 -6x - 4x^2 - 20x +24) (x^3 +5x^2 -6x - 4x^2 - 20x +24) reduces to x^3 +x^2 - 26x +24

OpenStudy (anonymous):

Wow! Thank you soo much! :))

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

do you understand where i got (x+6) * (x-1) * (x-4) from?

OpenStudy (anonymous):

yes, from the roots of the polynomial in the question.